OpenStudy (anonymous):

In the triangle below, calculate the value of x. A. 10 B. 5 C. 4.75 D. 11.85

OpenStudy (anonymous):

ques...?

OpenStudy (anonymous):

sorry i have been trying to post the picture but openstudy is being extremely slow .

OpenStudy (campbell_st):

you will need to use the sin ratio.. does that make sense..?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

x/12=sin57*=>x=10.06=> answer A

OpenStudy (anonymous):

how do you get that ? could you show me what you did ?

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

\[\sin \theta =\] p/h sin57=x/12 x=10